Find the probability of selecting 5 math books and 4 history books from a stack that contains 6 math books and 4 history books if the books are selected at random.
Answer by stanbon(75887)

Find the probability of selecting 5 math books and 4 history books from a stack that contains 6 math books and 4 history books if the books are selected at random.
------------------
# of ways to succeed:
# of ways to select 5 math: 6C5 = 6
# of ways to select 4 hist: 4C4 = 1
----
# of ways to succeed: 6*1 = 6
----------------
# of ways to select 9 books randomly: 10C9 = 10
-----
P(selecting 5 math and 4 hist) = 6/10 = 3/5
============================================
Cheers,
Stan H.